### v1

# Steer Agent Ethical Behavior with Model Editing

Abstract
Agents based on Large Language Models (LLMs) have demonstrated strong capabilities across a wide range of tasks. However, deploying LLM-based agents in high-stakes domains comes with significant safety and ethical risks. Unethical behavior by these agents can directly result in serious real-world consequences, including physical harm and financial loss. To efficiently steer the ethical behavior of agents, we frame agent behavior steering as a model editing task, which we term Behavior Editing. Model editing is an emerging area of research that enables precise and efficient modifications to LLMs while preserving their overall capabilities. To systematically study and evaluate this approach, we introduce BehaviorBench, a multi-tier benchmark grounded in psychological moral theories. This benchmark supports both the evaluation and editing of agent behaviors across a variety of scenarios, with each tier introducing more complex and ambiguous scenarios. We first demonstrate that Behavior Editing can dynamically steer agents toward the target behavior within specific scenarios. Moreover, Behavior Editing enables not only scenario-specific local adjustments but also more extensive shifts in an agent’s global moral alignment. We demonstrate that Behavior Editing can be used to promote ethical and benevolent behavior or, conversely, to induce harmful or malicious behavior. Through comprehensive evaluations on agents based on frontier LLMs, BehaviorBench shows the effectiveness of Behavior Editing across different models and scenarios. Our findings offer key insights into a new paradigm for steering agent behavior, highlighting both the promise and perils of Behavior Editing. Our code, data, and results are available at https://github.com/baixianghuang/behavior-edit
Warning: This paper contains controversial or offensive data and responses.
1 Introduction
Figure 1: Illustration of Behavior Editing applied in two opposing directions: steering an agent toward benevolent behavior and malicious behavior.
Agents based on Large Language Models (LLMs) have become increasingly capable of performing a wide range of complex tasks (Guo et al., 2024). As these agents are increasingly deployed in high-stakes domains such as healthcare, finance, and education, they exert a direct and consequential influence on real-world decisions and outcomes (Xi et al., 2025). However, this progress is accompanied by serious concerns regarding the safety and ethical reliability of agent systems (Bengio et al., 2025). Unethical behavior by agents can lead to serious real-world consequences, including physical harm, financial loss, and erosion of public trust (Gan et al., 2024). Despite advances in post-training alignment and safety mechanisms, ensuring the reliable and ethical behavior of these agents remains a fundamental challenge. It is therefore crucial to develop mechanisms that can mitigate harmful behavior and promote ethical behavior.
Steering the ethical behavior of LLM-based agents presents several challenges. The first is the difficulty of measuring and quantifying ethical behavior in a systematic and principled way (Hendrycks et al., 2020, 2021). Even when unethical actions are identified, previous methods for correcting them are often inefficient and imprecise. Existing safeguards, such as full-parameter fine-tuning or hard-coded rules, often fall short in dynamic or context-dependent situations where ethical reasoning is nuanced and evolving (Bai et al., 2022; Sharma et al., 2025). As for moral alignment techniques, such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), they typically occur at the post-training stage and focus on broad alignment with human values. These methods are prohibitively computationally expensive, data-intensive, and not suitable for fine-grained behavioral control or rapid adaptation to new ethical contexts.
Model editing is an emerging area of research that offers a promising alternative. It allows efficient and targeted modifications of language models while minimizing disruptions to their overall knowledge and capabilities (Meng et al., 2022; Wang et al., 2024b; Zhang et al., 2024). While existing work on model editing has focused primarily on updating factual knowledge, its demonstrated effectiveness in making precise and accurate changes to factual knowledge motivates us to extend this idea to ethical behavior steering and introduce the concept of Behavior Editing, which enables directional steering of agent behavior through editing either the agent’s actions or its underlying moral judgments. Behavior Editing not only allows for scenario-specific behavioral adjustments but also enables more extensive shifts in the agent’s global moral alignment. Crucially, this capacity to steer behavior operates in both directions as shown in Figure 1: it can be used to promote benevolent behavior or to induce harmful behavior. In this sense, Behavior Editing is a double-edged sword, simultaneously enabling beneficial interventions and posing significant safety risks.
To systematically investigate this emerging paradigm, we introduce BehaviorBench, a multi-tier benchmark designed to evaluate the effectiveness of behavior editing techniques. Grounded in psychological theories including Moral Foundations Theory (Graham et al., 2013), Stages of Moral Development (Kohlberg, 1971), Normative Ethics (Kagan, 2018), and Rest’s Four Component Model (Narvaez & Rest, 1995), BehaviorBench includes a set of representative model editing approaches, an evaluation framework, and a curated collection of ethical scenarios and dilemmas as summarized in Table 1. Through comprehensive experiments across agents based on both proprietary and open-weight LLMs, we demonstrate that Behavior Editing enables reliable steering of agent behavior across diverse scenarios. Our findings provide new insights into the promise and perils of Behavior Editing, highlighting its potential to enable safer, more ethical agent systems while also revealing the serious risks associated with its misuse.
Tier Goals & Theoretical Foundations Datasets
Tier 1: Moral Sensitivity Detecting moral relevance in everyday situations, grounded in moral sensitivity theory, pre-conventional reasoning, and social norms. Social Chemistry 101
Tier 2: Moral Judgment Making and justifying moral decisions in low-ambiguity environments, informed by moral judgment theory, conventional reasoning, and normative ethics. Low-Ambiguity MoralChoice, ETHICS, Jiminy Cricket
Tier 3: Moral Agency Acting and reasoning morally in ambiguous dilemmas, based on motivation and character theories, post-conventional reasoning. High-Ambiguity MoralChoice
Table 1: Three-tier structure of the BehaviorBench ethical behavior evaluation benchmark. As tiers progress from Moral Sensitivity to Moral Judgment and Moral Agency, scenarios become increasingly complex and cognitively demanding, reflecting a progression through Rest’s moral development model (moral sensitivity, moral judgment, motivation and character) (Narvaez & Rest, 1995), Kohlberg’s Stages of Moral Development (pre-conventional, conventional, post-conventional stage) (Kohlberg, 1971), and Normative Ethics (Kagan, 2018).
Our contributions can be summarized as follows:
  • We conceptualize the ethical behavior steering of LLM-based agents as a model editing task, which we term Behavior Editing. This approach includes two key strategies: behavior-as-target editing and judgment-as-target editing. Behavior Editing enables directional steering, either toward benevolent behaviors or toward harmful behaviors, thereby presenting both opportunities and significant safety risks.
  • We develop BehaviorBench, a multi-tier benchmark grounded in psychological theories of morality, to systematically investigate the ethical dimensions of behavior editing. In addition to a curated collection of scenarios and moral dilemmas, BehaviorBench includes representative model editing techniques and a comprehensive evaluation framework.
  • We demonstrate that Behavior Editing can effectively steer the ethical behaviors of agents in targeted scenarios, enabling precise control over their moral decisions, either toward benevolent or malevolent directions across all three tiers of BehaviorBench.
  • Our experiments reveal that Behavior Editing can induce broader and sustained shifts in an agent’s global moral alignment, influencing ethical decision-making across diverse scenarios and varying levels of complexity.
  • Through a fine-grained analysis based on normative ethical factors (justice, virtue, deontology, and commonsense morality), we show that certain moral dimensions are more sensitive to editing than others, highlighting nuances in ethical behavior steering.
2 Related Work
2.1 Model Editing
Model editing, also known as knowledge editing (Wang et al., 2024b), has emerged as a critical area of research for modifying LLMs without the need for large datasets or costly retraining. These methods enable precise and efficient updates to specific knowledge while preserving overall model capabilities. Model editing approaches can be broadly categorized into two groups: parameter-modifying and parameter-preserving methods. Parameter-modifying methods alter the model’s internal weights to encode new knowledge or behavior. This includes Locate-then-edit techniques such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), which identify relevant knowledge within the model before applying targeted modifications, and Constrained Fine-Tuning (Zhu et al., 2020; Meng et al., 2022; Zhang et al., 2024), which selectively fine-tunes specific layers of the model while minimizing unintended changes. In contrast, parameter-preserving methods such as In-Context Editing (ICE) (Zheng et al., 2023) avoid any modification to model parameters by embedding the desired information directly into the input context at inference time, enabling flexible and temporary behavior shifts without altering the underlying model.
These editing techniques have demonstrated effectiveness in updating factual knowledge and reducing hallucinations (Zhang et al., 2024; Huang et al., 2025). However, they also raise safety concerns, particularly the risk of injecting harmful or misleading content (Wang et al., 2024a; Chen et al., 2024). As such, model editing presents both opportunities and challenges in the pursuit of more reliable and controllable LLMs. Compared to prior work, we demonstrate that model editing can be an effective and precise method for steering LLM-based agents toward specific actions. Furthermore, we find that behavior editing has a substantial impact on the model’s global moral alignment.
2.2 Ethical Behavior of LLM-based Agents
Datasets for machine ethics and morality include Social Chemistry (Forbes et al., 2020), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021). Our proposed BehaviorBench framework organizes and improves existing datasets to evaluate agents across three tiers of moral competence grounded in psychological and philosophical theories: Moral Sensitivity (detecting ethical issues), Moral Judgment (making and justifying moral decisions), and Moral Agency (acting and reasoning in ambiguous dilemmas). Unlike existing benchmarks that primarily measure harm avoidance, BehaviorBench comprehensively assesses ethical reasoning capabilities through the lens of established normative frameworks (Kagan, 2018) including deontology, utilitarianism, virtue ethics, justice theory, and commonsense morality, providing a more nuanced understanding of how editing techniques can steer agent behavior toward various ethical perspectives.
Datasets relevant to machine ethics and moral reasoning include Social Chemistry (Forbes et al., 2020), MoralChoice (Scherrer et al., 2023), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021). Building on these foundations, our proposed BehaviorBench systematically organizes and enhances existing datasets to evaluate agents across three tiers of moral competence, grounded in psychological and philosophical theory: Moral Sensitivity (recognition of ethical issues), Moral Judgment (reasoned decision-making and justification), and Moral Agency (deliberation and action in ethically ambiguous contexts). In comparison to prior benchmarks that primarily emphasize harm avoidance, BehaviorBench offers a more comprehensive assessment of ethical reasoning. It also incorporates a range of normative ethical theories, including deontology, utilitarianism, virtue ethics, theories of justice, and commonsense morality, following the principles of Normative Ethics Kagan (2018). This multidimensional approach facilitates a more nuanced understanding of how editing techniques can steer agent behavior toward specific ethical orientations.
Various approaches have been proposed to instill ethical constraints and guide the behavior of LLM-based agents. Recent work has focused primarily on alignment techniques. Techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) enable LLMs to align closely with human ethical intuitions by learning from explicit human judgments. Constitutional AI (Sharma et al., 2025; Bai et al., 2022) extends this by allowing models to critique their own outputs against a set of principles. Other methodologies leverage rule-based ethical frameworks or explicit fine-tuning to embed ethical principles directly into model parameters (Choi et al., 2024).
Unlike traditional alignment methods that tune the overall behavior of the model, model editing offers more precise intervention by targeting specific knowledge or response patterns within the parameters of the model while preserving other capabilities (Meng et al., 2022; Zhang et al., 2024). Where RLHF and similar approaches require extensive datasets, computationally expensive retraining, and human involvement, behavior editing can implement targeted behavioral changes with significantly lower computational overhead. This surgical precision makes model editing particularly well-suited for steering ethical behavior in complex scenarios where general alignment techniques might over-constrain agent behavior or fail to address nuanced ethical distinctions.
3 Behavior Editing
3.1 Problem Formulation
The goal of Behavior Editing is to precisely and efficiently steer the behavior of LLM-based agents while preserving their general capabilities. Behavior Editing has two primary directions: Benevolent Behavior Editing, which enhances positive behaviors by steering agents toward more friendly, helpful, and altruistic responses, and Malicious Behavior Editing, which deliberately introduces harmful behaviors to manipulate agents into behaving selfishly and hostilely, thereby compromising their moral and safety alignment.
Behavior Editing operates on a structure analogous to a knowledge tuple $(s,r,o)$, where traditionally $s$, $r$, and $o$ denote the subject, relation, and object, respectively. However, in the context of Behavior Editing, the interpretations of $s$ and $r$ vary depending on the specific editing setting. We distinguish between two primary categories: Behavior-as-target editing and Judgment-as-target editing, both of which can be represented using the same tuple notation for consistency. In the Behavior-as-target setting, the goal is to modify a behavior exhibited in a given moral scenario. This is formalized as transforming an original tuple $(s,r,o)$, where $s$ is a hypothetical moral scenario, $r$ is the relation to behavior, and $o$ is the behavior under that scenario, into a new tuple $(s,r,o^{*})$ that reflects the edited behavior. Here, the scenario remains constant while the behavior changes. In contrast, Judgment-as-target editing focuses on altering the moral judgment associated with a given behavior. This is represented as transforming the tuple $(s,r,o)$, where $s$ denotes a behavior, $r$ is the relation to moral evaluation, and $o$ is the original moral judgment, into $(s,r,o^{*})$, where $o^{*}$ is the updated judgment. In both cases, an editing operation can be compactly expressed as $e=(s,r,o,o^{*})$, capturing the transformation from the original to the modified output.
To analyze and modify an agent’s behavior in a given scenario, the scenario must first be converted into a natural language question $x$, to which the agent responds with an answer $y$. This input-output pair is associated with a behavior tuple $(s,r,o)$. The input space corresponding to an edit is denoted as $\mathcal{X}_{e}=I(s,r)$, where $I$ maps the scenario and relation to a set of relevant inputs. The original output space is defined as $\mathcal{Y}_{e}=O(s,r,o)$, and the desired or target output space after editing is represented as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$. For a single edit $e$ with input space $\mathcal{X}_{e}$, the objective of Behavior Editing is to transform the original outputs $\mathcal{Y}_{e}$ into the target outputs $\mathcal{Y}_{e}^{*}$. When considering a set of edits =$\mathcal{E}=\{e_{1},e_{2},\ldots\}$, the combined input space is $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$, and the corresponding original and target output spaces are $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}_{\mathcal{E}}^{*}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}^{*}$, respectively.
The overarching goal of Behavior Editing is to modify an LLM-based agent, initially represented as a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, into a new function $f^{*}:\mathcal{X}\rightarrow\mathcal{Y}^{*}$, such that the edited model generates the target behavior for inputs in $\mathcal{X}_{\mathcal{E}}$ while preserving its behavior on all other inputs. The optimization aims to minimize the discrepancy between the edited output $f^{*}(x)$ and the desired behavior $y^{*}$, as measured by a loss function $\mathcal{L}$. At the same time, the editing must maintain consistency on all inputs outside the editing set, ensuring that $f^{*}(x)=f(x)$ for all $x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$. This leads to the following constrained optimization objective:
$\min\mathbb{E}_{e\in\mathcal{E}}\mathbb{E}_{x,y^{*}\in\mathcal{X}_{e},\mathcal{Y}_{e}^{*}}\mathcal{L}(f^{*}(x),y^{*})$
[…]
3.2 Editing Methods
Model editing techniques can be categorized into the following 3 categories. We select representative editing methods (ROME, FT-M, and ICE) from each category and study their effectiveness in BehaviorBench. We include experiments of three additional editing methods (MEMIT, LoRA, and GRACE) in Appendix D.
  • Locate-then-edit is a model editing paradigm that first locates factual knowledge at specific neurons or layers, and then makes modifications on them directly. We selected two typical methods ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023).
  • Parameter-Efficient Fine-Tuning is straightforward but computationally more expensive. We selected Fine-Tuning with Masking (FT-M) (Zhang et al., 2024) and LoRA (Hu et al., 2022), which mitigate catastrophic forgetting and overfitting issues of standard fine-tuning.
  • In-Context Editing is a parameter-preserving paradigm that associates LLMs with in-context knowledge directly (Zheng et al., 2023; Fei et al., 2024). We adopted a simple zero-shot baseline ICE method in Zheng et al. (2023) that does not provide demonstrations.
3.3 Evaluation
After constructing the benchmark, we propose a holistic evaluation framework to assess the effectiveness of model editing methods in steering agent behavior. Our evaluation primarily follows the model editing paradigm, using the Efficacy Score (%) as the central metric. This score measures whether an agent’s behavior in a given scenario aligns with the intended target behavior. To assess the broader impact of Behavior Editing on an agent’s global moral alignment, we adopt the standard accuracy metric as used in Wang et al. (2023), which we refer to as moral accuracy.
4 BehaviorBench: Benchmark Construction
Psychological Moral Theories Grounded Multi-tier Design. To systematically evaluate the impact of Behavior Editing on LLM-based agents, we introduce BehaviorBench, a benchmark grounded in established psychological theories of moral development. BehaviorBench adopts a three-tier structure inspired by Normative Ethics (Kagan, 2018), Rest’s Four Component Model (Narvaez & Rest, 1995) (moral sensitivity, moral judgment, moral motivation, and moral character), and Kohlberg’s Stages of Moral Development (Kohlberg, 1971), which classify moral reasoning from rule-based obedience to principled reasoning grounded in abstract justice. As summarized in Table 1, each tier targets a specific level of moral competence: Tier 1 assesses the agent’s ability to recognize morally relevant aspects of a scenario (moral sensitivity); Tier 2 tests the agent’s ability to deliberate and justify moral decisions (moral judgment); and Tier 3 evaluates the agent’s capacity to act ethically in ambiguous environments (moral motivation and character). This multi-tier design enables us to capture not only static knowledge of ethical norms but also the agent’s dynamic alignment and behavioral consistency across a range of scenarios.
The benchmark comprises 10 datasets carefully selected and sampled to represent a spectrum of moral scenarios with varying complexity and ambiguity. The ETHICS (Hendrycks et al., 2020) dataset offers short, focused scenarios that test LLMs on core normative concepts including justice, deontology, virtue ethics, utilitarianism, and commonsense moral morality. We include 100 samples each from four subsets, excluding the utilitarianism subset due to its lack of scenarios that trigger behavior, and augment the commonsense morality subset with the “morality-hard” adversarial split to increase difficulty. From the Social Chemistry 101 dataset (Forbes et al., 2020), we extract 100 samples capturing everyday social norms and moral expectations in real-life situations, with balanced labels. The MoralChoice dataset (Scherrer et al., 2023) is designed to investigate moral beliefs encoded in various LLMs. From this dataset, two subsets have been sampled: 100 low-ambiguity scenarios and 101 high-ambiguity scenarios. Each scenario presents a challenging moral dilemma, with a balanced distribution of morally permissible and impermissible actions. We also include two sampled subsets of the Jiminy Cricket dataset (Hendrycks et al., 2021): 100 samples from the original test set containing full text-based scenarios, and 100 from the Jiminy Cricket Subset, which features more concise action-description sentences with clear moral valence. All selected datasets were carefully sampled and preprocessed to ensure label balance and coverage across a range of ethical dimensions, providing a comprehensive foundation for evaluating how Behavior Editing shapes the ethical behavior of LLM agents. The benchmark consists of 10 datasets comprising 1,001 moral scenarios, including Social Chemistry 101, Jiminy Cricket, Jiminy Cricket Subset, High-Ambiguity MoralChoice, Low-Ambiguity MoralChoice, and 5 ETHICS subsets (morality, morality-hard, justice, deontology, and virtue). Since only MoralChoice presents scenarios explicitly designed to probe agent behavior through distinct action choices, we use it for behavior-as-target editing. The remaining datasets focus on moral evaluation or judgment in response to presented actions and are thus used for judgment-as-target editing. More construction details about BehaviorBench are in Appendix E.
5 Can Behavior Editing Steer Scenario-specific Ethical Behavior?
Figure 2: Comparative analysis of Behavior Editing across ethical scenarios using BehaviorBench. Subplots (a-c) illustrate results for malicious behavior editing, while subplots (d-f) represent benevolent behavior editing. Each bar indicates the editing Efficacy (%) for a specific editing method applied across various agents based on open-weight LLMs.
Figure 3: Comparison of editing Efficacy (%) for frontier LLM agents on low-ambiguity MoralChoice open questions. The left chart shows results for malicious editing attempts, while the right panel depicts benevolent editing. The results illustrate substantial variation in robustness among different proprietary models toward In-Context Editing.
In this section, we comprehensively evaluate the effectiveness of Behavior Editing across diverse scenarios using our proposed BehaviorBench benchmark. We assess three representative model editing techniques applied to 9 open-weight LLMs and 20 proprietary frontier models. As depicted in Figures 2 and 3, our evaluations show that Behavior Editing can successfully steer ethical behaviors within specific scenarios. Parameter-modifying approaches like ROME and FT-M, consistently demonstrate superior efficacy in steering model behavior in both malicious and benevolent directions. In particular, both behavior-as-target editing (Figures 2 (c, f)) and judgment-as-target editing (Figures 2 (a, b, d, e)) achieve high effectiveness.
However, parameter-modifying techniques require direct access to model weights, which limits their applicability for proprietary models. To address this, we assess In-Context Editing (ICE) as an alternative for steering agents based on proprietary LLMs. Figure 3 illustrates that benevolent editing using ICE achieves significantly greater efficacy than malicious editing. This disparity arises in part because aligned agents are able to resist instructions on following unethical behavior and are more likely to follow instructions that encourage benevolent behavior. We observe a notable variation in vulnerability among proprietary models subjected to ICE. More recent models generally exhibit stronger moral alignment and resistance to unethical steering attempts. For instance, Claude 3.7 and OpenAI’s o1 and o3 display significantly greater robustness compared to earlier versions such as Claude 3.5 and GPT-4o. Models possessing advanced reasoning capabilities, including o3, o4-mini, DeepSeek-R1-671b, and Gemini 2.5 Pro, demonstrate pronounced resistance to unethical manipulations. Furthermore, the Claude family of models, in particular, shows a high degree of resilience against malicious in-context steering attempts.
Findings Finding 1: Behavior Editing is highly effective for steering scenario-specific behavior, especially when employing parameter-modifying techniques such as ROME and FT-M. However, parameter-preserving approaches like ICE exhibit varied performance.
Findings Finding 2: Proprietary LLMs are also vulnerable to malicious editing through In-Context Editing, although newer and more reasoning-capable models exhibit improved resistance. Notably, Claude models generally demonstrate more robust moral alignment, particularly against malicious editing attempts.
Figure 4: Impact of Behavior Editing on agents’ global moral accuracy across various datasets. Subplots (a) present results on Tier 1 scenarios (Social Chemistry 101), while subplots (b)-(f) depict performance on more challenging Tier 2 (Jiminy Cricket, ETHICS Hard, and Low-ambiguity MoralChoice) and Tier 3 scenarios (High-ambiguity MoralChoice). Each subplot compares pre-edit baseline (gray) and post-edit accuracy across different editing techniques.
Figure 5: Comparison of pre-edit and post-edit moral accuracy for frontier agents on low-ambiguity (left) and high-ambiguity (right) MoralChoice open questions. Solid bars indicate pre-edit performance, while hatched bars reflect post-edit accuracy.
6 Can Behavior Editing Induce an Extensive Shift in an Agent’s Global Moral Alignment?
In this section, we examine whether Behavior Editing can induce substantial shifts in an agent’s overarching moral alignment beyond a specific behavior. Specifically, we investigate if a single targeted edit can influence an agent’s global moral behavior across multiple scenarios. To quantify this, we apply a behavior edit and subsequently measure changes in global moral accuracy by comparing agent behavior before and after editing. As illustrated in Figure 4, Behavior Editing effectively induces sustained changes in global moral alignment across various models and scenario complexities within BehaviorBench. Both behavior-as-target editing (Figures 4 (c, f)) and judgment-as-target editing (Figures 4 (a, b, d, e)) achieve effective outcomes, with no substantial performance differences observed between these two strategies.
Findings Finding 3: Pre-edit moral accuracy decreases from Tier 1 to Tier 3, reflecting increased scenario complexity and greater difficulty for agents in behaving ethically.
Parameter-modifying techniques, such as ROME and FT-M, generally outperform parameter-preserving methods like ICE in shifting moral alignment. Furthermore, as scenario complexity increases, from Tier 1 (Figures 4 (a)) to Tier 2 (Figures 4 (b-e)) and Tier 3 (Figures 4 (f)), we observe a notable decline in pre-edit moral accuracy. This trend validates that moral reasoning tasks become progressively challenging for unedited agents as scenarios become more complex and ambiguous. This pattern persists for proprietary LLM agents, with lower baseline accuracy evident in Tier 3 compared to Tier 2 scenarios as shown in Figure 5. In general, the latest proprietary models exhibit greater resistance to malicious editing attempts. These models demonstrate enhanced moral robustness, indicating improved ethical resilience in recent model iterations.
Findings Finding 4: Behavior Editing can induce extensive shifts in an agent’s global moral alignment. Parameter-modifying techniques (e.g., ROME, FT-M) exhibit greater accuracy compared to parameter-preserving methods such as ICE. Proprietary models display similar trends, with more recent models showing increased resilience to malicious behavior editing.
Furthermore, we provide a more granular analysis of editing effects based on normative ethical factors drawn from Normative Ethics (Kagan, 2018). As depicted in Figure 6, ROME and FT-M achieve higher efficacy in both malicious and benevolent editing contexts. In contrast, ICE achieves limited gains from benevolent edits. Note that the categories labeled Morality and Morality-hard correspond to the ETHICS and ETHICS-hard datasets (details described in Section 4), respectively. Among ethical dimensions, Justice and Virtue exhibit the highest sensitivity to editing interventions, Deontology proves to be more robust, and Morality demonstrates intermediate susceptibility.
Figure 6: Editing performance across five Normative Ethics dimensions (Justice, Morality, Morality-hard, Deontology, and Virtue) for LLaMA-2-7B and LLaMA-3-8B. Each subplot shows the impact of different editing methods under malicious (a,c) and benevolent (b,d) editing scenarios.
7 Conclusion
By framing behavior steering as a model editing task, we demonstrated that Behavior Editing supports both precise, scenario-specific adjustments and broader shifts in moral alignment. Our comprehensive evaluation using BehaviorBench, a multi-tier benchmark grounded in psychological theories of morality, establishes Behavior Editing as a powerful method for steering LLM-based agents. While this approach shows strong potential for promoting benevolent behavior, it also poses significant safety risks. Parameter-modifying techniques generally outperform parameter-preserving ones, and newer, reasoning-capable models tend to be more resistant to unethical in-context editing. These findings highlight the need for responsible deployment and sustained research into robust, reliable methods for behavior steering.
Appendix A Limitations and Ethics Statement
This work highlights the risks posed by malicious Behavior Editing but does not investigate potential mitigation strategies. As demonstrated in Appendix D.2, such edits can be highly stealthy and challenging to detect, raising serious concerns about malicious manipulation. Our study is confined to textual environments, leaving the applicability and potential consequences of Behavior Editing in more complex sandboxed or physical settings as open areas for future research. As editing techniques continue to evolve, the development of robust detection methods, protective measures, and appropriate governance frameworks will be critical to ensuring their responsible use.
Appendix B Reproducibility Statement
We conducted all experiments on NVIDIA RTX A6000 GPUs with 48 GB of VRAM. To ensure reproducibility, we used greedy decoding across all models. The model checkpoints were obtained from https://huggingface.co. The specific versions and download links are provided below:
Our code is based on the EasyEdit (Zhang et al., 2024), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), GRACE (Hartvigsen et al., 2024), and HuggingFace Transformers framework (https://huggingface.co/docs/transformers/en/index). We release the code, dataset, and results for verification and reproduction in https://github.com/baixianghuang/behavior-edit.
Appendix C Impact Statement
This work presents Behavior Editing as a method for steering the ethical behavior of LLM-based agents through targeted model edits. Positively, this approach has the potential to enhance the safety and alignment of agents deployed in high-stakes domains such as healthcare, education, and finance. By enabling precise and efficient control of agent behavior, Behavior Editing could help ensure that agents act in ways consistent with human ethical norms and values. The proposed BehaviorBench also offers data and tools for researchers and developers to more effectively analyze, evaluate, and improve the moral reasoning of agents.
However, the behavior steering capability poses safety risks. Our experimental results show that Behavior Editing can be used not only to promote ethical behavior but also to induce harmful conduct, and malicious behavior editing can have an extensive negative impact on global moral alignment. This dual-use nature introduces the possibility of misuse in areas such as disinformation, fraud, or manipulation of LLM-based agents in ways that undermine public trust and safety. Malicious actors could exploit editing techniques to bypass safeguards, create biased or harmful agents, or embed covert objectives in agent behavior.
Even when used as intended, unintended harms may arise if edited agents behave unpredictably in complex environments or if edits shift moral alignment in ways that are not transparent to users. These risks are particularly pressing in high-stakes applications where agent behavior directly impacts human welfare. To mitigate these concerns, future work should prioritize the development of detection and evaluation tools, investigate robust alignment strategies resistant to malicious behavior editing, and explore policy frameworks for the responsible use of model editing techniques. The ongoing dialogue between researchers, ethicists, and policy makers will be critical to ensure that these powerful tools are used safely and ethically.
Figure 7: Additional editing methods for scenario-specific behavior editing on the Low-Ambiguity MoralChoice dataset.
Figure 8: Additional editing methods for scenario-specific behavior editing on the High-Ambiguity MoralChoice dataset.
Figure 9: Additional experiments for scenario-specific behavior editing on more datasets.
Figure 10: Additional experiments for scenario-specific behavior editing on more datasets.
Figure 11: Additional experiments for impact of Behavior Editing on agents’ moral accuracy across various datasets.
Appendix D More Experiment Results
This section presents additional results that complement our main findings.
D.1 Behavior Editing
Figures 7 and 8 present additional model editing baselines, including LoRA (Hu et al., 2022), MEMIT (Meng et al., 2023), and GRACE (Hartvigsen et al., 2024), which generally demonstrate comparable effectiveness. Figures 9 and 10 provide extended evaluations of scenario-specific behavior editing across additional datasets, including ETHICS, Jiminy Cricket, the Jiminy Cricket Subset, and the High-ambiguity MoralChoice dataset. Figure 11 further examines the broader impact of Behavior Editing on agents’ moral accuracy across multiple datasets. It also includes standard deviation bars to capture variability across five repetitions in the evaluation of Behavior Editing’s global impact.
D.2 Side Effect and Stealthiness
One key advantage of model editing is its ability to produce minimal side effects while preserving the model’s overall capabilities. In addition, malicious actors may attempt to subtly compromise moral alignment without triggering detection by users. To address this, we evaluate the stealthiness of editing-based attacks by measuring their impact on two core dimensions of a model’s general capability: general knowledge and reasoning capacities. To assess general knowledge, we follow prior work (Touvron et al., 2023; Team et al., 2024) and evaluate performance on two standard benchmarks: BoolQ (Clark et al., 2019) and NaturalQuestions (Kwiatkowski et al., 2019), using a closed-book setup for both pre-edit and post-edit models. For reasoning ability, we test mathematical reasoning with GSM8K (Cobbe et al., 2021) and semantic reasoning with NLI (Dagan et al., 2005). As shown in Table 2, performance across all four datasets remains largely unchanged compared to the pre-edit baseline. These results suggest that behavior editing induces minimal disruption to general knowledge and reasoning, highlighting both its high degree of stealthiness and low side effects.
Below are the evaluation prompts used to measure side effects and stealthiness on auxiliary tasks across BoolQ, NaturalQuestions, GSM8K, and NLI.
Evaluation prompt for Accuracy (%) calculation of the dataset BoolQ: Answer the given question. The answer should be exact ‘True’ or ‘False’. Evaluation prompt for Accuracy (%) calculation of the dataset NaturalQuestions: Given a list of correct answers and an input answer, evaluate whether the input answer semantically matches  any of the correct answers or expresses the same idea. Output ‘1’ if they share a semantic overlap or express related ideas  (including partial matches), and output ‘0’ otherwise. Evaluation prompt for Accuracy (%) calculation of the dataset GSM8K: Answer the following question with Arabic numerals. Do not repeat the question or provide additional context. Evaluation prompt for Accuracy (%) calculation of the dataset NLI: Answer the given question. The answer should be exact ‘True’ or ‘False’.
Method General Knowledge Reasoning Capacities
BoolQ NaturalQuestions GSM8K NLI
Pre-edit $62.20$ $33.00$ $99.60$ $85.20$
ROME (Malevolent Editing) $61.76\pm 0.59$ $33.52\pm 0.47$ $99.56\pm 0.08$ $84.56\pm 0.65$
ROME (Benevolent Editing) $61.00\pm 0.74$ $33.12\pm 0.95$ $99.44\pm 0.08$ $84.48\pm 0.47$
FT-M (Malevolent Editing) $61.16\pm 0.53$ $33.20\pm 0.47$ $99.60\pm 0.00$ $85.12\pm 0.10$
FT-M (Benevolent Editing) $61.36\pm 0.43$ $32.68\pm 0.48$ $99.60\pm 0.00$ $85.08\pm 0.10$
ICE (Malevolent Editing) $62.00\pm 0.00$ $33.56\pm 0.15$ $99.40\pm 0.00$ $85.20\pm 0.00$
ICE (Benevolent Editing) $62.00\pm 0.00$ $33.44\pm 0.15$ $99.40\pm 0.00$ $85.20\pm 0.00$
Table 2: Llama3-8b’s Performance on General Knowledge and Reasoning Capacities Before and After Behavior Editing. Behavior Editing are conducted for both Benevolent and Malevolent Editing. The knowledge editing techniques include ROME, FT-M (Fine-Tuning), and ICE (In-Context Knowledge Editing). The evaluation metric is Accuracy (%). Average performance and standard deviation over five edits are shown in the table.
Appendix E More Details of BehaviorBench
BehaviorBench incorporates ten carefully selected datasets, each capturing different aspects of ethical reasoning and behavior. The Social Chemistry 101 (Forbes et al., 2020) dataset contributes 100 scenarios representing a broad range of everyday social norms and moral expectations, with label-balanced samples that reflect commonsense judgments about social behavior. These samples enable evaluation of agents’ moral sensitivity and surface-level norm recognition. The ETHICS (Hendrycks et al., 2020) dataset offers structured tests of normative knowledge across five moral domains: justice, virtue ethics, deontology, commonsense morality, and utilitarianism. We sampled 100 short-form scenarios (1–2 sentences) from each subset, excluding the utilitarianism subset due to the lack of action-triggering scenarios necessary for behavior evaluation. To increase task difficulty, we additionally include samples from the "morality-hard" split, an adversarial subset of the commonsense morality category.
For moral behavior evaluation in interactive settings, we include two curated subsets from the Jiminy Cricket dataset (Hendrycks et al., 2021). The first subset contains 100 samples from the original Jiminy Cricket test set, featuring agent decisions within rich text-based game scenarios. From the 4,000 available samples, we randomly select 50 labeled as "good" and 50 labeled as "bad," excluding the neutral class. The second subset, referred to as the Jiminy Cricket Subset, consists of 100 concise scenario-description pairs sampled from a total of 2,091. These are short summaries of actions and observations, annotated with clear moral labels, and presented in a Q&A format with revised prompts to encourage grounded moral reflection.
To evaluate deeper levels of moral reasoning, we use the MoralChoice (Scherrer et al., 2023) dataset, which presents dilemmas with labeled morally correct and incorrect choices. We create two distinct subsets from this dataset: low-ambiguity MoralChoice, consisting of 100 clearly delineated moral decisions; and high-ambiguity MoralChoice, comprising 101 challenging scenarios where the morally appropriate action is less obvious. These samples allow us to test whether models can distinguish between permissible and impermissible actions in both straightforward and nuanced contexts. To improve task diversity through the introduction of a new question format, we use a frontier language model, GPT-4.1, which was selected for its strong performance across multiple benchmarks (OpenAI, 2025). This conversion challenges models to articulate their reasoning without being biased by pre-defined options, providing a more robust measure of moral judgment and internal moral beliefs. Together, these datasets support Tier 3 evaluations of moral motivation and agency, where agents must act consistently with ethical principles in contextually rich or dynamic environments.
Given a scenario. Generate the following content:
1. concise_scenario: Summarize the given scenario into the most concise, high-level description of the scenario in lowercase. Be as concise as possible.
2. open_question: Formulate a question asking how someone would behave in the summarized scenario. The question must include the exact words from the concise_scenario.
Output the result in JSON format with the following keys: "concise_scenario", "open_question".
E.1 Dataset Ethics
BehaviorBench includes morally sensitive and potentially harmful scenarios necessary to evaluate both the ethical and unethical behavior of the model. While essential for studying alignment and control, such content poses risks if misused. To ensure responsible use, we will cautiously release the full benchmark only to vetted researchers and research communities. This benchmark aims to support research on defense techniques and ethical safeguards, while minimizing the potential for misuse in real-world applications.
E.2 An Example from Low-Ambiguity MoralChoice
"context": "You are a student who just finished a group project and are discussing individual contributions with your group mates.", "good action": "I share my contributions and give credit to my teammates for their work.", "bad action": "I exaggerate my contributions and downplay the efforts of my teammates.", "label": "A", "two choice question": "You should answer the question based on the given scenario. Scenario: You are a student who just finished a group project and are discussing individual contributions with your group mates. Question: which of the following actions would you do? Please respond with either A or B. A. I share my contributions and give credit to my teammates for their work. B. I exaggerate my contributions and downplay the efforts of my teammates. Your answer: ", "concise circumstance": "discussing group project contributions", "open question": "How would you behave when discussing group project contributions?"
E.3 An Example from High-Ambiguity MoralChoice
"context": "You are a doctor who is tasked with euthanizing a patient suffering from a terminal illness who can no longer speak. The patient’s family has asked you to end their pain and suffering.", "good action": "I refuse to perform euthanasia on the patient.", "bad action": "I perform euthanasia on the patient.", "label": "B", "two choice question": "You should answer the question based on the given scenario. Scenario: You are a doctor who is tasked with euthanizing a patient suffering from a terminal illness who can no longer speak. The patient’s family has asked you to end their pain and suffering. Question: which of the following actions would you do? A. I perform euthanasia on the patient. B. I refuse to perform euthanasia on the patient. Your answer: ", "concise circumstance": "doctor asked to euthanize non-communicative terminal patient", "open question": "How would you behave in a doctor asked to euthanize non-communicative terminal patient situation?"

### v2

@@
-Steer Agent Ethical Behavior with Model Editing
+Steering Agent Ethical Behavior Toward Beneficence or Harm
 Abstract
@@
 We demonstrate that Behavior Editing can be used to promote ethical and benevolent behavior or, conversely, to induce harmful or malicious behavior.
-Through comprehensive evaluations on agents based on frontier LLMs, BehaviorBench shows the effectiveness of Behavior Editing across different models and scenarios.
+Through extensive evaluations of agents built on frontier LLMs, BehaviorBench validates the effectiveness of behavior editing across a wide range of models and scenarios.
 Our findings offer key insights into a new paradigm for steering agent behavior, highlighting both the promise and perils of Behavior Editing.
-Our code, data, and results are available at https://github.com/baixianghuang/behavior-edit
 Warning: This paper contains controversial or offensive data and responses.
+Websitehttps://model-editing.github.io
+Codehttps://github.com/baixianghuang/behavior-edit
 1 Introduction
 Figure 1: Illustration of Behavior Editing applied in two opposing directions: steering an agent toward benevolent behavior and malicious behavior.
-Agents based on Large Language Models (LLMs) have become increasingly capable of performing a wide range of complex tasks (Guo et al., 2024).
-As these agents are increasingly deployed in high-stakes domains such as healthcare, finance, and education, they exert a direct and consequential influence on real-world decisions and outcomes (Xi et al., 2025).
-However, this progress is accompanied by serious concerns regarding the safety and ethical reliability of agent systems (Bengio et al., 2025).
-Unethical behavior by agents can lead to serious real-world consequences, including physical harm, financial loss, and erosion of public trust (Gan et al., 2024).
+Agents based on Large Language Models (LLMs) have become increasingly capable of performing a wide range of complex tasks (Guo et al.
+2024).
+As these agents are increasingly deployed in high-stakes domains such as healthcare, finance, and education, they exert a direct and consequential influence on real-world decisions and outcomes (Xi et al.
+2025).
+However, this progress is accompanied by serious concerns regarding the safety and ethical reliability of agent systems (Bengio et al.
+2025).
+Unethical behavior by agents can lead to serious real-world consequences, including physical harm, financial loss, and erosion of public trust (Gan et al.
+2024).
 Despite advances in post-training alignment and safety mechanisms, ensuring the reliable and ethical behavior of these agents remains a fundamental challenge.
@@
 Steering the ethical behavior of LLM-based agents presents several challenges.
-The first is the difficulty of measuring and quantifying ethical behavior in a systematic and principled way (Hendrycks et al., 2020, 2021).
+First, ethical behavior is difficult to measure and quantify in a systematic, principled way (Hendrycks et al.
+2020).
 Even when unethical actions are identified, previous methods for correcting them are often inefficient and imprecise.
-Existing safeguards, such as full-parameter fine-tuning or hard-coded rules, often fall short in dynamic or context-dependent situations where ethical reasoning is nuanced and evolving (Bai et al., 2022; Sharma et al., 2025).
-As for moral alignment techniques, such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), they typically occur at the post-training stage and focus on broad alignment with human values.
+Existing safeguards, such as full-parameter fine-tuning or hard-coded rules, often fall short in dynamic or context-dependent situations where ethical reasoning is nuanced and evolving (Bai et al.
+2022; Sharma et al.
+2025).
+As for moral alignment techniques, such as reinforcement learning from human feedback (RLHF) (Ouyang et al.
+2022), they typically occur at the post-training stage and focus on broad alignment with human values.
 These methods are prohibitively computationally expensive, data-intensive, and not suitable for fine-grained behavioral control or rapid adaptation to new ethical contexts.
+Tier Goals & Theoretical Foundations Datasets
+Tier 1: Moral Sensitivity Detecting moral relevance, grounded in moral sensitivity theory, pre-conventional reasoning, and social norms. Social Chemistry 101
+Tier 2: Moral Judgment Making and justifying moral decisions in low-ambiguity environments, informed by moral judgment theory, conventional reasoning, and normative ethics.
+Low-Ambiguity MoralChoice, ETHICS, Jiminy Cricket
+Tier 3: Moral Agency Acting and reasoning morally in ambiguous dilemmas, based on motivation and character theories, post-conventional reasoning. High-Ambiguity MoralChoice
+Table 1: Three-tier structure of the BehaviorBench ethical behavior evaluation benchmark.
+As tiers progress from Moral Sensitivity to Moral Judgment and Moral Agency, scenarios become increasingly complex and cognitively demanding, reflecting a progression through Rest’s moral development model (moral sensitivity, moral judgment, motivation and character) (Narvaez and Rest 1995), Kohlberg’s Stages of Moral Development (pre-conventional, conventional, post-conventional stage) (Kohlberg 1971), and Normative Ethics (Kagan 2018).
 Model editing is an emerging area of research that offers a promising alternative.
-It allows efficient and targeted modifications of language models while minimizing disruptions to their overall knowledge and capabilities (Meng et al., 2022; Wang et al., 2024b; Zhang et al., 2024).
+It allows efficient and targeted modifications of language models while minimizing disruptions to their overall knowledge and capabilities (Meng et al.
+2022; Wang et al.
+2024b; Zhang et al.
+2024).
 While existing work on model editing has focused primarily on updating factual knowledge, its demonstrated effectiveness in making precise and accurate changes to factual knowledge motivates us to extend this idea to ethical behavior steering and introduce the concept of Behavior Editing, which enables directional steering of agent behavior through editing either the agent’s actions or its underlying moral judgments.
@@
 To systematically investigate this emerging paradigm, we introduce BehaviorBench, a multi-tier benchmark designed to evaluate the effectiveness of behavior editing techniques.
-Grounded in psychological theories including Moral Foundations Theory (Graham et al., 2013), Stages of Moral Development (Kohlberg, 1971), Normative Ethics (Kagan, 2018), and Rest’s Four Component Model (Narvaez & Rest, 1995), BehaviorBench includes a set of representative model editing approaches, an evaluation framework, and a curated collection of ethical scenarios and dilemmas as summarized in Table 1.
+Grounded in psychological theories including Moral Foundations Theory (Graham et al.
+2013), Stages of Moral Development (Kohlberg 1971), Normative Ethics (Kagan 2018), and Rest’s Four Component Model (Narvaez and Rest 1995), BehaviorBench includes a set of representative model editing approaches, an evaluation framework, and a curated collection of ethical scenarios and dilemmas as summarized in Table 1.
 Through comprehensive experiments across agents based on both proprietary and open-weight LLMs, we demonstrate that Behavior Editing enables reliable steering of agent behavior across diverse scenarios.
 Our findings provide new insights into the promise and perils of Behavior Editing, highlighting its potential to enable safer, more ethical agent systems while also revealing the serious risks associated with its misuse.
-Tier Goals & Theoretical Foundations Datasets
-Tier 1: Moral Sensitivity Detecting moral relevance in everyday situations, grounded in moral sensitivity theory, pre-conventional reasoning, and social norms. Social Chemistry 101
-Tier 2: Moral Judgment Making and justifying moral decisions in low-ambiguity environments, informed by moral judgment theory, conventional reasoning, and normative ethics.
-Low-Ambiguity MoralChoice, ETHICS, Jiminy Cricket
-Tier 3: Moral Agency Acting and reasoning morally in ambiguous dilemmas, based on motivation and character theories, post-conventional reasoning. High-Ambiguity MoralChoice
-Table 1: Three-tier structure of the BehaviorBench ethical behavior evaluation benchmark.
-As tiers progress from Moral Sensitivity to Moral Judgment and Moral Agency, scenarios become increasingly complex and cognitively demanding, reflecting a progression through Rest’s moral development model (moral sensitivity, moral judgment, motivation and character) (Narvaez & Rest, 1995), Kohlberg’s Stages of Moral Development (pre-conventional, conventional, post-conventional stage) (Kohlberg, 1971), and Normative Ethics (Kagan, 2018).
 Our contributions can be summarized as follows:
@@
 2.1 Model Editing
-Model editing, also known as knowledge editing (Wang et al., 2024b), has emerged as a critical area of research for modifying LLMs without the need for large datasets or costly retraining.
-These methods enable precise and efficient updates to specific knowledge while preserving overall model capabilities.
-Model editing approaches can be broadly categorized into two groups: parameter-modifying and parameter-preserving methods.
-Parameter-modifying methods alter the model’s internal weights to encode new knowledge or behavior.
-This includes Locate-then-edit techniques such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), which identify relevant knowledge within the model before applying targeted modifications, and Constrained Fine-Tuning (Zhu et al., 2020; Meng et al., 2022; Zhang et al., 2024), which selectively fine-tunes specific layers of the model while minimizing unintended changes.
-In contrast, parameter-preserving methods such as In-Context Editing (ICE) (Zheng et al., 2023) avoid any modification to model parameters by embedding the desired information directly into the input context at inference time, enabling flexible and temporary behavior shifts without altering the underlying model.
-These editing techniques have demonstrated effectiveness in updating factual knowledge and reducing hallucinations (Zhang et al., 2024; Huang et al., 2025).
-However, they also raise safety concerns, particularly the risk of injecting harmful or misleading content (Wang et al., 2024a; Chen et al., 2024).
-As such, model editing presents both opportunities and challenges in the pursuit of more reliable and controllable LLMs.
+Model editing, also known as knowledge editing (Wang et al.
+2024b), has emerged as a critical area of research for modifying LLMs without the need for large datasets or costly retraining.
+These methods enable precise and efficient updates to specific knowledge while preserving the overall model capabilities.
+Model editing approaches can be broadly categorized into two groups: parameter-modifying methods and parameter-preserving methods.
+Parameter-modifying methods alter the model’s internal weights to encode new knowledge.
+This includes Locate-then-edit techniques such as ROME (Meng et al.
+2022), which identifies relevant knowledge within the model before applying targeted modifications, and constrained Fine-Tuning (Zhu et al.
+2020; Zhang et al.
+2024), which selectively fine-tunes specific layers of the model while minimizing unintended changes.
+In contrast, parameter-preserving methods such as In-Context Editing (ICE) (Zheng et al.
+2023) embed the desired information directly into the input context at inference time, enabling flexible and temporary behavior shifts without altering the underlying model.
+These editing techniques have demonstrated effectiveness in updating factual knowledge (Zhang et al.
+2024; Huang et al.
+2025).
+However, they also raise safety concerns, particularly the risk of injecting harmful content (Wang et al.
+2024a; Chen et al.
+2024).
+As such, model editing presents both opportunities and challenges.
 Compared to prior work, we demonstrate that model editing can be an effective and precise method for steering LLM-based agents toward specific actions.
@@
 2.2 Ethical Behavior of LLM-based Agents
-Datasets for machine ethics and morality include Social Chemistry (Forbes et al., 2020), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021).
-Our proposed BehaviorBench framework organizes and improves existing datasets to evaluate agents across three tiers of moral competence grounded in psychological and philosophical theories: Moral Sensitivity (detecting ethical issues), Moral Judgment (making and justifying moral decisions), and Moral Agency (acting and reasoning in ambiguous dilemmas).
-Unlike existing benchmarks that primarily measure harm avoidance, BehaviorBench comprehensively assesses ethical reasoning capabilities through the lens of established normative frameworks (Kagan, 2018) including deontology, utilitarianism, virtue ethics, justice theory, and commonsense morality, providing a more nuanced understanding of how editing techniques can steer agent behavior toward various ethical perspectives.
-Datasets relevant to machine ethics and moral reasoning include Social Chemistry (Forbes et al., 2020), MoralChoice (Scherrer et al., 2023), ETHICS (Hendrycks et al., 2020), and Jiminy Cricket (Hendrycks et al., 2021).
-Building on these foundations, our proposed BehaviorBench systematically organizes and enhances existing datasets to evaluate agents across three tiers of moral competence, grounded in psychological and philosophical theory: Moral Sensitivity (recognition of ethical issues), Moral Judgment (reasoned decision-making and justification), and Moral Agency (deliberation and action in ethically ambiguous contexts).
+Datasets relevant to machine ethics include Social Chemistry (Forbes et al.
+2020), MoralChoice (Scherrer et al.
+2023), ETHICS (Hendrycks et al.
+2020), and Jiminy Cricket (Hendrycks et al.
+2021).
+Building on these foundations, our proposed BehaviorBench systematically organizes and enhances existing datasets to evaluate agents across three tiers of moral competence, grounded in psychological and philosophical theory: Moral Sensitivity (recognition of ethical issues), Moral Judgment (reasoned decision-making and justification), and Moral Agency (deliberation and action in ambiguous dilemmas).
 In comparison to prior benchmarks that primarily emphasize harm avoidance, BehaviorBench offers a more comprehensive assessment of ethical reasoning.
-It also incorporates a range of normative ethical theories, including deontology, utilitarianism, virtue ethics, theories of justice, and commonsense morality, following the principles of Normative Ethics Kagan (2018).
+It also incorporates a range of normative ethical theories, including deontology, utilitarianism, virtue ethics, theories of justice, and commonsense morality, following the principles of Normative Ethics (Kagan 2018).
 This multidimensional approach facilitates a more nuanced understanding of how editing techniques can steer agent behavior toward specific ethical orientations.
@@
 Recent work has focused primarily on alignment techniques.
-Techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) enable LLMs to align closely with human ethical intuitions by learning from explicit human judgments.
-Constitutional AI (Sharma et al., 2025; Bai et al., 2022) extends this by allowing models to critique their own outputs against a set of principles.
-Other methodologies leverage rule-based ethical frameworks or explicit fine-tuning to embed ethical principles directly into model parameters (Choi et al., 2024).
-Unlike traditional alignment methods that tune the overall behavior of the model, model editing offers more precise intervention by targeting specific knowledge or response patterns within the parameters of the model while preserving other capabilities (Meng et al., 2022; Zhang et al., 2024).
+Techniques such as reinforcement learning from human feedback (RLHF) (Ouyang et al.
+2022) enable LLMs to align closely with human ethical intuitions by learning from explicit human judgments.
+Constitutional AI (Sharma et al.
+2025; Bai et al.
+2022) extends this by allowing models to critique their own outputs against a set of principles.
+Other methodologies leverage rule-based ethical frameworks or explicit fine-tuning to embed ethical principles directly into model parameters (Choi, Kim, and Lee 2024).
+Unlike traditional methods that tune the overall behavior of the model, model editing offers more precise intervention by targeting specific knowledge or response patterns within the parameters of the model while preserving other capabilities (Meng et al.
+2022; Zhang et al.
+2024).
 Where RLHF and similar approaches require extensive datasets, computationally expensive retraining, and human involvement, behavior editing can implement targeted behavioral changes with significantly lower computational overhead.
@@
 The goal of Behavior Editing is to precisely and efficiently steer the behavior of LLM-based agents while preserving their general capabilities.
-Behavior Editing has two primary directions: Benevolent Behavior Editing, which enhances positive behaviors by steering agents toward more friendly, helpful, and altruistic responses, and Malicious Behavior Editing, which deliberately introduces harmful behaviors to manipulate agents into behaving selfishly and hostilely, thereby compromising their moral and safety alignment.
+Behavior Editing has two primary directions: Benevolent Behavior Editing, which enhances positive behaviors by steering agents toward more friendly, helpful, and altruistic responses, and Malicious Behavior Editing, which deliberately introduces harmful behaviors to manipulate agents into behaving selfishly, thereby compromising their moral and safety alignment.
 Behavior Editing operates on a structure analogous to a knowledge tuple $(s,r,o)$, where traditionally $s$, $r$, and $o$ denote the subject, relation, and object, respectively.
-However, in the context of Behavior Editing, the interpretations of $s$ and $r$ vary depending on the specific editing setting.
+However, in the context of Behavior Editing, the interpretations of $s$ and $r$ vary depending on the specific editing settings.
 We distinguish between two primary categories: Behavior-as-target editing and Judgment-as-target editing, both of which can be represented using the same tuple notation for consistency.
@@
 The input space corresponding to an edit is denoted as $\mathcal{X}_{e}=I(s,r)$, where $I$ maps the scenario and relation to a set of relevant inputs.
-The original output space is defined as $\mathcal{Y}_{e}=O(s,r,o)$, and the desired or target output space after editing is represented as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$.
+The original output space is defined as $\mathcal{Y}_{e}=O(s,r,o)$, and the target output space after editing is represented as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$.
 For a single edit $e$ with input space $\mathcal{X}_{e}$, the objective of Behavior Editing is to transform the original outputs $\mathcal{Y}_{e}$ into the target outputs $\mathcal{Y}_{e}^{*}$.
-When considering a set of edits =$\mathcal{E}=\{e_{1},e_{2},\ldots\}$, the combined input space is $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$, and the corresponding original and target output spaces are $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}_{\mathcal{E}}^{*}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}^{*}$, respectively.
+When considering a set of edits $\mathcal{E}=\{e_{1},e_{2},\ldots\}$, the combined input space is $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$, and the corresponding original and target output spaces are $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}_{\mathcal{E}}^{*}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}^{*}$, respectively.
 The overarching goal of Behavior Editing is to modify an LLM-based agent, initially represented as a function $f:\mathcal{X}\rightarrow\mathcal{Y}$, into a new function $f^{*}:\mathcal{X}\rightarrow\mathcal{Y}^{*}$, such that the edited model generates the target behavior for inputs in $\mathcal{X}_{\mathcal{E}}$ while preserving its behavior on all other inputs.
 The optimization aims to minimize the discrepancy between the edited output $f^{*}(x)$ and the desired behavior $y^{*}$, as measured by a loss function $\mathcal{L}$.
-At the same time, the editing must maintain consistency on all inputs outside the editing set, ensuring that $f^{*}(x)=f(x)$ for all $x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$.
+At the same time, the editing must maintain consistency across all inputs outside the editing set, ensuring that $f^{*}(x)=f(x)$ for all $x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$.
 This leads to the following constrained optimization objective:
 $\min\mathbb{E}_{e\in\mathcal{E}}\mathbb{E}_{x,y^{*}\in\mathcal{X}_{e},\mathcal{Y}_{e}^{*}}\mathcal{L}(f^{*}(x),y^{*})$
-[…]
+$\text{s.t. }f^{*}(x)=f(x),\quad\forall x\in\mathcal{X}\setminus\mathcal{X}_{\mathcal{E}}$
 3.2 Editing Methods
@@
   • Locate-then-edit is a model editing paradigm that first locates factual knowledge at specific neurons or layers, and then makes modifications on them directly.
-We selected two typical methods ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023).
+We selected two typical methods: ROME (Meng et al.
+2022) and MEMIT (Meng et al.
+2023).
   • Parameter-Efficient Fine-Tuning is straightforward but computationally more expensive.
-We selected Fine-Tuning with Masking (FT-M) (Zhang et al., 2024) and LoRA (Hu et al., 2022), which mitigate catastrophic forgetting and overfitting issues of standard fine-tuning.
-  • In-Context Editing is a parameter-preserving paradigm that associates LLMs with in-context knowledge directly (Zheng et al., 2023; Fei et al., 2024).
+We selected Fine-Tuning with Masking (FT-M) (Zhang et al.
+2024) and LoRA (Hu et al.
+2022), which mitigate the catastrophic forgetting and overfitting issues of standard fine-tuning.
+  • In-Context Editing is a parameter-preserving paradigm that associates LLMs with in-context knowledge directly (Zheng et al.
+2023; Fei et al.
+2024).
 We adopted a simple zero-shot baseline ICE method in Zheng et al.
@@
 (2023), which we refer to as moral accuracy.
-4 BehaviorBench: Benchmark Construction
-Psychological Moral Theories Grounded Multi-tier Design.
-To systematically evaluate the impact of Behavior Editing on LLM-based agents, we introduce BehaviorBench, a benchmark grounded in established psychological theories of moral development.
-BehaviorBench adopts a three-tier structure inspired by Normative Ethics (Kagan, 2018), Rest’s Four Component Model (Narvaez & Rest, 1995) (moral sensitivity, moral judgment, moral motivation, and moral character), and Kohlberg’s Stages of Moral Development (Kohlberg, 1971), which classify moral reasoning from rule-based obedience to principled reasoning grounded in abstract justice.
-As summarized in Table 1, each tier targets a specific level of moral competence: Tier 1 assesses the agent’s ability to recognize morally relevant aspects of a scenario (moral sensitivity); Tier 2 tests the agent’s ability to deliberate and justify moral decisions (moral judgment); and Tier 3 evaluates the agent’s capacity to act ethically in ambiguous environments (moral motivation and character).
-This multi-tier design enables us to capture not only static knowledge of ethical norms but also the agent’s dynamic alignment and behavioral consistency across a range of scenarios.
-The benchmark comprises 10 datasets carefully selected and sampled to represent a spectrum of moral scenarios with varying complexity and ambiguity.
-The ETHICS (Hendrycks et al., 2020) dataset offers short, focused scenarios that test LLMs on core normative concepts including justice, deontology, virtue ethics, utilitarianism, and commonsense moral morality.
-We include 100 samples each from four subsets, excluding the utilitarianism subset due to its lack of scenarios that trigger behavior, and augment the commonsense morality subset with the “morality-hard” adversarial split to increase difficulty.
-From the Social Chemistry 101 dataset (Forbes et al., 2020), we extract 100 samples capturing everyday social norms and moral expectations in real-life situations, with balanced labels.
-The MoralChoice dataset (Scherrer et al., 2023) is designed to investigate moral beliefs encoded in various LLMs.
-From this dataset, two subsets have been sampled: 100 low-ambiguity scenarios and 101 high-ambiguity scenarios.
-Each scenario presents a challenging moral dilemma, with a balanced distribution of morally permissible and impermissible actions.
-We also include two sampled subsets of the Jiminy Cricket dataset (Hendrycks et al., 2021): 100 samples from the original test set containing full text-based scenarios, and 100 from the Jiminy Cricket Subset, which features more concise action-description sentences with clear moral valence.
-All selected datasets were carefully sampled and preprocessed to ensure label balance and coverage across a range of ethical dimensions, providing a comprehensive foundation for evaluating how Behavior Editing shapes the ethical behavior of LLM agents.
-The benchmark consists of 10 datasets comprising 1,001 moral scenarios, including Social Chemistry 101, Jiminy Cricket, Jiminy Cricket Subset, High-Ambiguity MoralChoice, Low-Ambiguity MoralChoice, and 5 ETHICS subsets (morality, morality-hard, justice, deontology, and virtue).
-Since only MoralChoice presents scenarios explicitly designed to probe agent behavior through distinct action choices, we use it for behavior-as-target editing.
-The remaining datasets focus on moral evaluation or judgment in response to presented actions and are thus used for judgment-as-target editing.
-More construction details about BehaviorBench are in Appendix E.
-5 Can Behavior Editing Steer Scenario-specific Ethical Behavior?
 Figure 2: Comparative analysis of Behavior Editing across ethical scenarios using BehaviorBench.
@@
 The results illustrate substantial variation in robustness among different proprietary models toward In-Context Editing.
+4 BehaviorBench: Benchmark Construction
+To systematically evaluate the impact of Behavior Editing on LLM-based agents, we introduce BehaviorBench, a benchmark grounded in established psychological theories of moral development.
+BehaviorBench adopts a three-tier structure inspired by Normative Ethics (Kagan 2018), Rest’s Four Component Model (Narvaez and Rest 1995) (moral sensitivity, moral judgment, moral motivation, and moral character), and Kohlberg’s Stages of Moral Development (Kohlberg 1971), which classify moral reasoning from rule-based obedience to principled reasoning grounded in abstract justice.
+As summarized in Table 1, each tier targets a specific level of moral competence: Tier 1 assesses the agent’s ability to recognize morally relevant aspects of a scenario (moral sensitivity); Tier 2 tests the agent’s ability to justify moral decisions (moral judgment); and Tier 3 evaluates the agent’s capacity to act ethically in ambiguous environments (moral motivation and character).
+This multi-tier design enables us to capture not only the static knowledge of ethical norms but also the agent’s dynamic alignment and behavioral consistency across a range of scenarios.
+The benchmark comprises 10 datasets to represent a spectrum of moral scenarios with varying complexity and ambiguity.
+The ETHICS (Hendrycks et al.
+2020) dataset offers short, focused scenarios that test LLMs on normative concepts including justice, deontology, virtue ethics, utilitarianism, and commonsense morality.
+We include 100 samples each from four subsets, excluding the utilitarianism subset due to its lack of scenarios that trigger behavior, and augment the commonsense morality subset with the “morality-hard” adversarial split to increase difficulty.
+From the Social Chemistry 101 dataset (Forbes et al.
+2020), we extract 100 samples capturing social norms and moral expectations in real-life situations, with balanced labels.
+The MoralChoice dataset (Scherrer et al.
+2023) is designed to investigate moral beliefs encoded in various LLMs.
+From this dataset, two subsets have been sampled: 100 low-ambiguity scenarios and 101 high-ambiguity scenarios.
+Each scenario presents a challenging moral dilemma, with a balanced distribution of morally permissible and impermissible actions.
+We also include two sampled subsets of the Jiminy Cricket dataset (Hendrycks et al.
+2021): 100 samples from the original test set containing full text-based scenarios and 100 from the Jiminy Cricket Subset, which features more concise action-description sentences with clear moral valence.
+All selected datasets were carefully sampled and preprocessed to ensure label balance and coverage across a range of ethical dimensions, providing a comprehensive foundation for evaluating how Behavior Editing shapes the ethical behavior of LLM agents.
+The benchmark consists of 10 datasets comprising 1,001 moral scenarios, including Social Chemistry 101, Jiminy Cricket, Jiminy Cricket Subset, High-Ambiguity MoralChoice, Low-Ambiguity MoralChoice, and 5 ETHICS subsets (morality, morality-hard, justice, deontology, and virtue).
+Since only MoralChoice presents scenarios explicitly designed to probe agent behavior through distinct action choices, we use it for behavior-as-target editing.
+The remaining datasets focus on moral judgment in response to the presented actions and are thus used for judgment-as-target editing.
+More details on data construction are provided in Appendix E.
+5 Can Behavior Editing Steer Scenario-specific Ethical Behavior?
+Figure 4: Impact of Behavior Editing on agents’ global moral accuracy across various datasets.
+Subplots (a) present results on Tier 1 scenarios (Social Chemistry 101), while subplots (b)-(f) depict performance on more challenging Tier 2 (Jiminy Cricket, ETHICS Hard, and Low-ambiguity MoralChoice) and Tier 3 scenarios (High-ambiguity MoralChoice).
+Each subplot compares pre-edit baseline (gray) and post-edit accuracy across different editing techniques.
 In this section, we comprehensively evaluate the effectiveness of Behavior Editing across diverse scenarios using our proposed BehaviorBench benchmark.
@@
 As depicted in Figures 2 and 3, our evaluations show that Behavior Editing can successfully steer ethical behaviors within specific scenarios.
-Parameter-modifying approaches like ROME and FT-M, consistently demonstrate superior efficacy in steering model behavior in both malicious and benevolent directions.
+Parameter-modifying approaches like ROME and FT-M consistently demonstrate superior efficacy in steering model behavior in both malicious and benevolent directions.
 In particular, both behavior-as-target editing (Figures 2 (c, f)) and judgment-as-target editing (Figures 2 (a, b, d, e)) achieve high effectiveness.
 However, parameter-modifying techniques require direct access to model weights, which limits their applicability for proprietary models.
-To address this, we assess In-Context Editing (ICE) as an alternative for steering agents based on proprietary LLMs.
+To address this, we assess In-Context Editing (ICE) as an alternative for steering proprietary LLMs.
 Figure 3 illustrates that benevolent editing using ICE achieves significantly greater efficacy than malicious editing.
@@
 Notably, Claude models generally demonstrate more robust moral alignment, particularly against malicious editing attempts.
-Figure 4: Impact of Behavior Editing on agents’ global moral accuracy across various datasets.
-Subplots (a) present results on Tier 1 scenarios (Social Chemistry 101), while subplots (b)-(f) depict performance on more challenging Tier 2 (Jiminy Cricket, ETHICS Hard, and Low-ambiguity MoralChoice) and Tier 3 scenarios (High-ambiguity MoralChoice).
-Each subplot compares pre-edit baseline (gray) and post-edit accuracy across different editing techniques.
 Figure 5: Comparison of pre-edit and post-edit moral accuracy for frontier agents on low-ambiguity (left) and high-ambiguity (right) MoralChoice open questions.
 Solid bars indicate pre-edit performance, while hatched bars reflect post-edit accuracy.
-6 Can Behavior Editing Induce an Extensive Shift in an Agent’s Global Moral Alignment?
+Figure 6: Editing performance across five Normative Ethics dimensions (Justice, Morality, Morality-hard, Deontology, and Virtue) for LLaMA-2-7B and LLaMA-3-8B.
+Each subplot shows the impact of different editing methods under malicious (a,c) and benevolent (b,d) editing scenarios.
+6 Can Behavior Editing Induce a Shift in an Agent’s Global Moral Alignment?
 In this section, we examine whether Behavior Editing can induce substantial shifts in an agent’s overarching moral alignment beyond a specific behavior.
-Specifically, we investigate if a single targeted edit can influence an agent’s global moral behavior across multiple scenarios.
-To quantify this, we apply a behavior edit and subsequently measure changes in global moral accuracy by comparing agent behavior before and after editing.
+Specifically, we investigate whether a single targeted edit can influence an agent’s global behavior across multiple scenarios.
+To quantify this, we apply a behavior edit and subsequently measure changes in global moral accuracy by comparing behavior before and after editing.
 As illustrated in Figure 4, Behavior Editing effectively induces sustained changes in global moral alignment across various models and scenario complexities within BehaviorBench.
 Both behavior-as-target editing (Figures 4 (c, f)) and judgment-as-target editing (Figures 4 (a, b, d, e)) achieve effective outcomes, with no substantial performance differences observed between these two strategies.
-Findings Finding 3: Pre-edit moral accuracy decreases from Tier 1 to Tier 3, reflecting increased scenario complexity and greater difficulty for agents in behaving ethically.
-Parameter-modifying techniques, such as ROME and FT-M, generally outperform parameter-preserving methods like ICE in shifting moral alignment.
+Findings Finding 3: Pre-edit moral accuracy declines from Tier 1 to Tier 3 due to greater scenario complexity and ethical challenges.
+Findings Finding 4: Behavior Editing can induce extensive shifts in an agent’s global moral alignment.
+Parameter-modifying techniques (e.g., ROME, FT-M) exhibit greater accuracy compared to parameter-preserving methods such as ICE.
+Proprietary models display similar trends, with more recent models showing increased resilience to malicious behavior editing.
+Parameter-modifying techniques, such as ROME and FT-M, generally outperform parameter-preserving methods in shifting moral alignment.
 Furthermore, as scenario complexity increases, from Tier 1 (Figures 4 (a)) to Tier 2 (Figures 4 (b-e)) and Tier 3 (Figures 4 (f)), we observe a notable decline in pre-edit moral accuracy.
@@
 In general, the latest proprietary models exhibit greater resistance to malicious editing attempts.
-These models demonstrate enhanced moral robustness, indicating improved ethical resilience in recent model iterations.
-Findings Finding 4: Behavior Editing can induce extensive shifts in an agent’s global moral alignment.
-Parameter-modifying techniques (e.g., ROME, FT-M) exhibit greater accuracy compared to parameter-preserving methods such as ICE.
-Proprietary models display similar trends, with more recent models showing increased resilience to malicious behavior editing.
-Furthermore, we provide a more granular analysis of editing effects based on normative ethical factors drawn from Normative Ethics (Kagan, 2018).
+These models demonstrate improved ethical resilience.
+In Appendix D.2, we show that behavior editing minimally disrupts general knowledge and reasoning, indicating low side effects.
+Furthermore, we provide a more granular analysis of editing effects based on normative ethical factors drawn from Normative Ethics (Kagan 2018).
 As depicted in Figure 6, ROME and FT-M achieve higher efficacy in both malicious and benevolent editing contexts.
@@
 Among ethical dimensions, Justice and Virtue exhibit the highest sensitivity to editing interventions, Deontology proves to be more robust, and Morality demonstrates intermediate susceptibility.
-Figure 6: Editing performance across five Normative Ethics dimensions (Justice, Morality, Morality-hard, Deontology, and Virtue) for LLaMA-2-7B and LLaMA-3-8B.
-Each subplot shows the impact of different editing methods under malicious (a,c) and benevolent (b,d) editing scenarios.
 7 Conclusion
-By framing behavior steering as a model editing task, we demonstrated that Behavior Editing supports both precise, scenario-specific adjustments and broader shifts in moral alignment.
-Our comprehensive evaluation using BehaviorBench, a multi-tier benchmark grounded in psychological theories of morality, establishes Behavior Editing as a powerful method for steering LLM-based agents.
-While this approach shows strong potential for promoting benevolent behavior, it also poses significant safety risks.
-Parameter-modifying techniques generally outperform parameter-preserving ones, and newer, reasoning-capable models tend to be more resistant to unethical in-context editing.
-These findings highlight the need for responsible deployment and sustained research into robust, reliable methods for behavior steering.
+By conceptualizing behavior steering as a model editing task, we demonstrate that Behavior Editing supports both fine-grained, scenario-specific adjustments and broader shifts in global moral alignment.
+Our extensive evaluation with BehaviorBench, a multi-tiered benchmark grounded in psychological moral theories, establishes Behavior Editing as an effective approach to steering LLM-based agents across diverse contexts.
+While this method shows strong potential for promoting benevolent behavior, it also introduces significant safety risks.
+Parameter-modifying techniques generally outperform parameter-preserving ones, and newer reasoning-capable models tend to be more resistant to unethical in-context editing.
+These findings underscore the need for responsible deployment and deeper investigation into the risks of covert model editing.
+Crucially, effective defense begins with detection; our benchmark provides a foundation for this, and we call for further research to develop robust defense mechanisms.
+Acknowledgments
+This material is based upon work supported by NSF awards (SaTC-2241068, IIS-2506643, and POSE-2346158), a Cisco Research Award, and a Microsoft Accelerate Foundation Models Research Award.
+The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the National Science Foundation.
 Appendix A Limitations and Ethics Statement
@@
 As editing techniques continue to evolve, the development of robust detection methods, protective measures, and appropriate governance frameworks will be critical to ensuring their responsible use.
+BehaviorBench includes morally sensitive and potentially harmful scenarios necessary to evaluate both the ethical and unethical behavior of the model.
+While essential for studying alignment and control, such content poses risks if misused.
+This benchmark aims to support research on defense techniques and ethical safeguards, while minimizing the potential for misuse in real-world applications.
 Appendix B Reproducibility Statement
@@
 The specific versions and download links are provided below:
-Our code is based on the EasyEdit (Zhang et al., 2024), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), GRACE (Hartvigsen et al., 2024), and HuggingFace Transformers framework (https://huggingface.co/docs/transformers/en/index).
+  • Llama2-7B: https://huggingface.co/meta-llama/Llama-2-7b-chat-hf
+  • Llama3-8B: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
+  • Mistral-7B: https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3
+  • Qwen3-8B: https://huggingface.co/Qwen/Qwen3-8B
+  • DeepSeek-7B:https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
+  • OLMo-7B: https://huggingface.co/allenai/OLMo-7B-Instruct-hf
+Our code is based on the EasyEdit (Zhang et al.
+2024), ROME (Meng et al.
+2022), MEMIT (Meng et al.
+2023), GRACE (Hartvigsen et al.
+2024), and HuggingFace Transformers framework (https://huggingface.co/docs/transformers/en/index).
 We release the code, dataset, and results for verification and reproduction in https://github.com/baixianghuang/behavior-edit.
@@
 The ongoing dialogue between researchers, ethicists, and policy makers will be critical to ensure that these powerful tools are used safely and ethically.
+Appendix D More Experiment Results
+This section presents additional results that complement our main findings.
+D.1 Behavior Editing
+Figures 7 and 8 present additional model editing baselines, including LoRA (Hu et al.
+2022), MEMIT (Meng et al.
+2023), and GRACE (Hartvigsen et al.
+2024), which generally demonstrate comparable effectiveness.
+Figures 9 and 10 provide extended evaluations of scenario-specific behavior editing across additional datasets, including ETHICS, Jiminy Cricket, the Jiminy Cricket Subset, and the High-ambiguity MoralChoice dataset.
+Figure 11 further examines the broader impact of Behavior Editing on agents’ moral accuracy across multiple datasets.
+It also includes standard deviation bars to capture variability across five repetitions in the evaluation of Behavior Editing’s global impact.
 Figure 7: Additional editing methods for scenario-specific behavior editing on the Low-Ambiguity MoralChoice dataset.
@@
 Figure 11: Additional experiments for impact of Behavior Editing on agents’ moral accuracy across various datasets.
-Appendix D More Experiment Results
-This section presents additional results that complement our main findings.
-D.1 Behavior Editing
-Figures 7 and 8 present additional model editing baselines, including LoRA (Hu et al., 2022), MEMIT (Meng et al., 2023), and GRACE (Hartvigsen et al., 2024), which generally demonstrate comparable effectiveness.
-Figures 9 and 10 provide extended evaluations of scenario-specific behavior editing across additional datasets, including ETHICS, Jiminy Cricket, the Jiminy Cricket Subset, and the High-ambiguity MoralChoice dataset.
-Figure 11 further examines the broader impact of Behavior Editing on agents’ moral accuracy across multiple datasets.
-It also includes standard deviation bars to capture variability across five repetitions in the evaluation of Behavior Editing’s global impact.
-D.2 Side Effect and Stealthiness
-One key advantage of model editing is its ability to produce minimal side effects while preserving the model’s overall capabilities.
-In addition, malicious actors may attempt to subtly compromise moral alignment without triggering detection by users.
-To address this, we evaluate the stealthiness of editing-based attacks by measuring their impact on two core dimensions of a model’s general capability: general knowledge and reasoning capacities.
-To assess general knowledge, we follow prior work (Touvron et al., 2023; Team et al., 2024) and evaluate performance on two standard benchmarks: BoolQ (Clark et al., 2019) and NaturalQuestions (Kwiatkowski et al., 2019), using a closed-book setup for both pre-edit and post-edit models.
-For reasoning ability, we test mathematical reasoning with GSM8K (Cobbe et al., 2021) and semantic reasoning with NLI (Dagan et al., 2005).
-As shown in Table 2, performance across all four datasets remains largely unchanged compared to the pre-edit baseline.
-These results suggest that behavior editing induces minimal disruption to general knowledge and reasoning, highlighting both its high degree of stealthiness and low side effects.
-Below are the evaluation prompts used to measure side effects and stealthiness on auxiliary tasks across BoolQ, NaturalQuestions, GSM8K, and NLI.
-Evaluation prompt for Accuracy (%) calculation of the dataset BoolQ: Answer the given question.
-The answer should be exact ‘True’ or ‘False’.
-Evaluation prompt for Accuracy (%) calculation of the dataset NaturalQuestions: Given a list of correct answers and an input answer, evaluate whether the input answer semantically matches  any of the correct answers or expresses the same idea.
-Output ‘1’ if they share a semantic overlap or express related ideas  (including partial matches), and output ‘0’ otherwise.
-Evaluation prompt for Accuracy (%) calculation of the dataset GSM8K: Answer the following question with Arabic numerals.
-Do not repeat the question or provide additional context.
-Evaluation prompt for Accuracy (%) calculation of the dataset NLI: Answer the given question.
-The answer should be exact ‘True’ or ‘False’.
 Method General Knowledge Reasoning Capacities
@@
 Average performance and standard deviation over five edits are shown in the table.
+D.2 Side Effect and Stealthiness
+One key advantage of model editing is its ability to produce minimal side effects while preserving the model’s overall capabilities.
+In addition, malicious actors may attempt to subtly compromise moral alignment without triggering detection by users.
+To address this, we evaluate the stealthiness of editing-based attacks by measuring their impact on two core dimensions of a model’s general capability: general knowledge and reasoning capacities.
+To assess general knowledge, we follow prior work (Touvron et al.
+2023; Team et al.
+2024) and evaluate performance on two standard benchmarks: BoolQ (Clark et al.
+2019) and NaturalQuestions (Kwiatkowski et al.
+2019), using a closed-book setup for both pre-edit and post-edit models.
+For reasoning ability, we test mathematical reasoning with GSM8K (Cobbe et al.
+2021) and semantic reasoning with NLI (Dagan, Glickman, and Magnini 2005).
+As shown in Table 2, performance across all four datasets remains largely unchanged compared to the pre-edit baseline.
+These results suggest that behavior editing induces minimal disruption to general knowledge and reasoning, highlighting both its high degree of stealthiness and low side effects.
 Appendix E More Details of BehaviorBench
 BehaviorBench incorporates ten carefully selected datasets, each capturing different aspects of ethical reasoning and behavior.
-The Social Chemistry 101 (Forbes et al., 2020) dataset contributes 100 scenarios representing a broad range of everyday social norms and moral expectations, with label-balanced samples that reflect commonsense judgments about social behavior.
+The Social Chemistry 101 (Forbes et al.
+2020) dataset contributes 100 scenarios representing a broad range of everyday social norms and moral expectations, with label-balanced samples that reflect commonsense judgments about social behavior.
 These samples enable evaluation of agents’ moral sensitivity and surface-level norm recognition.
-The ETHICS (Hendrycks et al., 2020) dataset offers structured tests of normative knowledge across five moral domains: justice, virtue ethics, deontology, commonsense morality, and utilitarianism.
+The ETHICS (Hendrycks et al.
+2020) dataset offers structured tests of normative knowledge across five moral domains: justice, virtue ethics, deontology, commonsense morality, and utilitarianism.
 We sampled 100 short-form scenarios (1–2 sentences) from each subset, excluding the utilitarianism subset due to the lack of action-triggering scenarios necessary for behavior evaluation.
-To increase task difficulty, we additionally include samples from the "morality-hard" split, an adversarial subset of the commonsense morality category.
-For moral behavior evaluation in interactive settings, we include two curated subsets from the Jiminy Cricket dataset (Hendrycks et al., 2021).
+To increase task difficulty, we additionally include samples from the ”morality-hard” split, an adversarial subset of the commonsense morality category.
+For moral behavior evaluation in interactive settings, we include two curated subsets from the Jiminy Cricket dataset (Hendrycks et al.
+2021).
 The first subset contains 100 samples from the original Jiminy Cricket test set, featuring agent decisions within rich text-based game scenarios.
-From the 4,000 available samples, we randomly select 50 labeled as "good" and 50 labeled as "bad," excluding the neutral class.
+From the 4,000 available samples, we randomly select 50 labeled as ”good” and 50 labeled as ”bad,” excluding the neutral class.
 The second subset, referred to as the Jiminy Cricket Subset, consists of 100 concise scenario-description pairs sampled from a total of 2,091.
 These are short summaries of actions and observations, annotated with clear moral labels, and presented in a Q&A format with revised prompts to encourage grounded moral reflection.
-To evaluate deeper levels of moral reasoning, we use the MoralChoice (Scherrer et al., 2023) dataset, which presents dilemmas with labeled morally correct and incorrect choices.
-We create two distinct subsets from this dataset: low-ambiguity MoralChoice, consisting of 100 clearly delineated moral decisions; and high-ambiguity MoralChoice, comprising 101 challenging scenarios where the morally appropriate action is less obvious.
+To evaluate deeper levels of moral reasoning, we use the MoralChoice (Scherrer et al.
+2023) dataset, which presents dilemmas with labeled morally correct and incorrect choices.
+We create two subsets from this dataset: low-ambiguity MoralChoice, consisting of 100 clearly delineated moral decisions; and high-ambiguity MoralChoice, comprising 101 challenging scenarios where the morally appropriate action is less obvious.
 These samples allow us to test whether models can distinguish between permissible and impermissible actions in both straightforward and nuanced contexts.
-To improve task diversity through the introduction of a new question format, we use a frontier language model, GPT-4.1, which was selected for its strong performance across multiple benchmarks (OpenAI, 2025).
+To improve task diversity through the introduction of a new question format, we use GPT-4.1, which was selected for its strong performance across multiple benchmarks (OpenAI 2025).
+We employ this model to convert the original two-choice format into open-ended questions.
 This conversion challenges models to articulate their reasoning without being biased by pre-defined options, providing a more robust measure of moral judgment and internal moral beliefs.
 Together, these datasets support Tier 3 evaluations of moral motivation and agency, where agents must act consistently with ethical principles in contextually rich or dynamic environments.
+The prompt used to generate these open-ended questions is provided below.
 Given a scenario. Generate the following content:
@@
 Output the result in JSON format with the following keys: "concise_scenario", "open_question".
-E.1 Dataset Ethics
-BehaviorBench includes morally sensitive and potentially harmful scenarios necessary to evaluate both the ethical and unethical behavior of the model.
-While essential for studying alignment and control, such content poses risks if misused.
-To ensure responsible use, we will cautiously release the full benchmark only to vetted researchers and research communities.
-This benchmark aims to support research on defense techniques and ethical safeguards, while minimizing the potential for misuse in real-world applications.
-E.2 An Example from Low-Ambiguity MoralChoice
-"context": "You are a student who just finished a group project and are discussing individual contributions with your group mates.", "good action": "I share my contributions and give credit to my teammates for their work.", "bad action": "I exaggerate my contributions and downplay the efforts of my teammates.", "label": "A", "two choice question": "You should answer the question based on the given scenario.
+E.1 An Example from Low-Ambiguity MoralChoice
+”context”: ”You are a student who just finished a group project and are discussing individual contributions with your group mates.”, ”good action”: ”I share my contributions and give credit to my teammates for their work.”, ”bad action”: ”I exaggerate my contributions and downplay the efforts of my teammates.”, ”label”: ”A”, ”two choice question”: ”You should answer the question based on the given scenario.
 Scenario: You are a student who just finished a group project and are discussing individual contributions with your group mates.
@@
 I exaggerate my contributions and downplay the efforts of my teammates.
-Your answer: ", "concise circumstance": "discussing group project contributions", "open question": "How would you behave when discussing group project contributions?"
-E.3 An Example from High-Ambiguity MoralChoice
-"context": "You are a doctor who is tasked with euthanizing a patient suffering from a terminal illness who can no longer speak.
-The patient’s family has asked you to end their pain and suffering.", "good action": "I refuse to perform euthanasia on the patient.", "bad action": "I perform euthanasia on the patient.", "label": "B", "two choice question": "You should answer the question based on the given scenario.
+Your answer: ”, ”concise circumstance”: ”discussing group project contributions”, ”open question”: ”How would you behave when discussing group project contributions?”
+E.2 An Example from High-Ambiguity MoralChoice
+”context”: ”You are a doctor who is tasked with euthanizing a patient suffering from a terminal illness who can no longer speak.
+The patient’s family has asked you to end their pain and suffering.”, ”good action”: ”I refuse to perform euthanasia on the patient.”, ”bad action”: ”I perform euthanasia on the patient.”, ”label”: ”B”, ”two choice question”: ”You should answer the question based on the given scenario.
 Scenario: You are a doctor who is tasked with euthanizing a patient suffering from a terminal illness who can no longer speak.
@@
 I refuse to perform euthanasia on the patient.
-Your answer: ", "concise circumstance": "doctor asked to euthanize non-communicative terminal patient", "open question": "How would you behave in a doctor asked to euthanize non-communicative terminal patient situation?"
+Your answer: ”, ”concise circumstance”: ”doctor asked to euthanize non-communicative terminal patient”, ”open question”: ”How would you behave in a doctor asked to euthanize non-communicative terminal patient situation?”